# FMGS: Foundation Model Embedded 3D Gaussian Splatting for Holistic 3D Scene Understanding

### v1

@@ 5.1 Object Detection in the Wild
Average 18.0% 54.8% 83.0% 93.2%
Table 1: Accuracy of object detection with open-vocabulary queries. comparison between Feature Fields Distillation [26] using LSeg [27] features (FFD-Lseg), OWL-ViT [36], LERF [25] and Ours FMGS. We highlight the best, second-best accuracy scores. Please find more details on scenes and text queries for LERF dataset in [25].
@@
Figure 7: Object detection results. Left: The Ground-truth bounding boxes (blue), our detected highest-relevancy pixel (green) and the one detected by LERF (red) [25]. Middle: Ours relevancy score with the given text query. The query word is shown at the most left of each row. Right: LERF’s relevancy score with the given text query. Our computed relevancy score is more focused on the target objects linked to the query.
@@
Table 2: Segmentation Evaluation. We report the mIoU($\uparrow$) scores and the mAP($\uparrow$) scores of the following methods in 6 scenes of 3D-OVS dataset [30]. Note that 3D-OVS is a weakly supervised method, which knows the segmentation annotations in training and specially designed for segmentation task. Our method and LERF are 3D method training without any segmentation annotations, relying only on the relevancy between class query and the rendered CLIP features. OV-Seg [28] is a supervised method for segmentation task. Our method and LERF are unsupervised method, under apple-to-apple comparison.
@@
Table 3: Ablation study. Object detection comparison between our full method, ours without dot product similarity (dotpsim) loss, and ours without hybrid CLIP features by averaging at multiple scales for supervision, using single scale Clip feature at the finest-resolution instead.
@@
Inference Runtime: Our FMGS, relying on 3D Gaussian Splatting rendering [24], excels in efficiently rendering RGB images. We’ve implemented our rendering method for CLIP and DINO feature maps based on a CUDA implementation of Gaussian Splatting rendering. Even when rendering deep features with high dimensions, which can significantly increase computation time, our FMGS remains remarkably fast. It can render the $480\times 270$ CLIP feature map, DINO feature map, and RGB image jointly at an impressively high rate of $103.4$ FPS during inference, even with our unoptimized implementation. In contrast, LERF operates at a significantly slower pace, achieving a mere $0.1214$ FPS during inference. This slowness stems from LERF’s need to perform a brute-force search for the best scales when rendering CLIP features, spanning a range from 0 to 2 meters with 30 increments. Consequently, we are 851.73 times faster than LERF in rendering CLIP features, enabling efficient real-time open-vocabulary queries after our scene representation is trained.
Figure 8: Semantic segmentation results. In the rows from top to bottom, we display RGB images, ground-truth (Gt) segmentation masks, our segmentation results, and the segmentation results obtained by LERF [25] scene representation. It’s essential to note that neither our method nor LERF was initially intended for the segmentation task. Our primary aim is to evaluate the pixel accuracy of the relevance map computed from the rendered CLIP features.
@@ 5.2 Unsupervised Segmentation
In following experiments we use FMGS to segment queries and evaluate their segmentation masks. Note that our method is not delicately designed for segmentation task. We lack a dedicated segmentation header for predicting segmentation masks, nor do we explicitly partition the scene at the object level. We have examined the open-vocabulary language understanding capability of FMGS in the object detection experiments discussed in the above section. Our primary objective for doing this segmentation evaluation is to assess the pixel-level accuracy of the rendered CLIP features obtained from the trained scene representation. Segmentation relies on matching these rendered CLIP features to the embeddings of the provided semantic labels.
@@
Baselines: We conduct a direct comparison of our method with LERF [25]. To perform semantic segmentation, we initially obtain relevancy scores by computing the cosine similarity between the rendered CLIP feature and the embeddings of all class labels (this is different from the relevancy score calculation with auxiliary canonical phrases involved in Sec. 4.3.). These relevancy scores serve as segmentation logits, and we subsequently apply the softmax function to convert them into probabilities. Each pixel is then assigned a semantic class label corresponding to the maximum probability. Note that LERF [25] requires a scale factor when rendering CLIP features, and we report the best segmentation results that can be achieved by LERF by selecting the best scales for each ray. It’s also important to note that both LERF and our method encounter challenges in discerning the semantic labels of backgrounds when presented with visibility-limited close views and lack of context. Therefore, we have replaced the original background labels, including ‘white sheet’, ‘wood wall’, ‘grey sofa,’ and ‘lime wall,’ with a more general label ‘background’ when testing LERF and our method.
Additionally, for comprehensive reference, we present results obtained using the dedicated 3D-OVS method [30] for the segmentation task. However, it is worth emphasizing that comparing object detection methods like ours and LERF [25] to 3D-OVS is not entirely equitable, as acknowledged in the paper of 3D-OVS [30]. 3D-OVS [30] has prior access to segmentation class labels and distill class-related information into the radiance field during training. In contrast, neither LERF nor our methods have access to class labels during scene representation training. Consequently, the trained 3D-OVS scene representation can only be effectively employed for querying the classes known before training, and does not support arbitrary semantic queries beyond the trained classes. Furthermore, we compare to a 2D ceiling approach [28], OV-Seg, which is directly trained for open-vocabulary semantic segmentation by fine-tuning CLIP on masked image regions and text descriptions. OV-Seg is supervised with mask-category pairs, while ours and LERF are purely unsupervised.
Evaluation Results The segmentation experiment results are presented in Table 2 and Figure 8. Notably, our approach outperforms LERF [25] by a significant margin across all cases. This superior performance can be attributed to the higher quality of our rendered CLIP feature compared to the one produced by LERF. Our method exhibits more concentrated high relevancy around the queried objects, showcasing the advantage of our semantic scene representation, which maintains high semantic consistency in local proximity.
@@ 5.3.1 Hybrid CLIP feature
In this ablation study, we investigated using a single scale of CLIP features, rather than our hybrid CLIP features, which are obtained by averaging multiple-scale CLIP features extracted from patches at different resolutions. As demonstrated in Table 3, the hybrid CLIP feature for supervision is greatly important. The scene understanding capability of our method is severely compromised when employing only a single-scale CLIP feature for supervision.
@@ 6 Discussion and Limitations
In terms of limitations, FMGS currently relies heavily on the presence of high-quality and calibrated input images, a limitation shared with NeRF-based approaches. Additionally, the performance of FMGS is entirely contingent on the quality of the base foundation models used for training the feature fields. It is conceivable that a model better suited for localizing language within images could yield improved feature field quality. Futhermore, for improved performance in the semantic segmentation task, it is advisable to embed a dedicated segmentation foundation model, such as SAM [9], into our scene representation. Alternatively, a straightforward approach for semantic segmentation is to initially segment the images using the foundation model and then assign semantic meanings to the segments based on our rendered CLIP features.

### v2

@@ 5.1 Object Detection in the Wild
Average Acc. 18.0% 54.8% 83.0% 93.2%
Inference FPS - - 0.1214 103.4
Table 1: Accuracy and runtime efficiency (Frames Per Second, FPS) of object detection with open-vocabulary queries. comparison between Feature Fields Distillation [29] using LSeg [30] features (FFD-Lseg), OWL-ViT [39], LERF [27] and Ours FMGS. We highlight the best, second-best accuracy scores. Please find more details on scenes and text queries for LERF dataset in [27].
@@
Figure 7: Object detection results. The left and right groups of figures illustrate the detection results for four open-vocabulary queries, respectively. Each group comprises three subfigures: Left displays the ground-truth bounding boxes (blue), our detected highest-relevancy pixel (green) and the one detected by LERF (red) [27]. Middle showcases our relevancy score corresponding to the given text query. The text query is shown at the far left of each row. Right showcases LERF’s relevancy score corresponding to the given text query. Our computed relevancy score is more focused on the target objects linked to the query.
@@
Ours-Refined 80.6 85.5 90.8 97.4 92.6 98.5 87.9 97.8 84.5 94.8 89.4 97.2
Table 2: Segmentation Evaluation. We report the mIoU($\uparrow$) scores and the mAP($\uparrow$) scores of the following methods in 6 scenes of 3D-OVS dataset [33]. Note that 3D-OVS is a weakly supervised method, which knows the segmentation annotations in training and specially designed for segmentation task. Our method and LERF are 3D method training without any segmentation annotations, relying only on the relevancy between class query and the rendered CLIP features. OV-Seg [31] is a supervised method for segmentation task. Our method and LERF are unsupervised methods, under apple-to-apple comparison. We can further post-process and refine our 2D segmentation results by SAM [28] and get the results shown as ‘Ours-Refined’.
@@
W/ LERF CLIP 91.7 70.5 72.5 72.5 87.0 78.8
W/O MHE 91.7 71.8 90.0 90.6 77.8 84.4
Table 3: Ablation study. Object detection comparison between our full method, ours without dot product similarity (dotpsim) loss, and ours without hybrid CLIP features by averaging at multiple scales for supervision, using single scale CLIP feature at the finest-resolution instead, as well as ours with LERF CLIP at multiple individual scales [27].
@@
Inference Runtime: Our FMGS, relying on 3D Gaussian Splatting rendering [26], excels in efficiently rendering RGB images. We’ve implemented our rendering method for CLIP and DINO feature maps based on a CUDA implementation of Gaussian Splatting rendering. Even when rendering deep features with high dimensions, which can significantly increase computation time, our FMGS remains remarkably fast. It can render the $480\times 270$ CLIP feature map, DINO feature map, and RGB image jointly at an impressively high rate of $103.4$ FPS during inference, even with our unoptimized implementation. In contrast, LERF operates at a significantly slower pace, achieving a mere $0.1214$ FPS during inference (see Table 1). This slowness stems from LERF’s need to perform a brute-force search for the best scales when rendering CLIP features, spanning a range from 0 to 2 meters with 30 increments. Consequently, we are 851.73 times faster than LERF in rendering CLIP features, enabling efficient real-time open-vocabulary queries after our scene representation is trained.
Figure 8: Semantic segmentation results. In the rows from top to bottom, we display RGB images, ground-truth (GT) segmentation masks, our refined segmentation results, our segmentation results, and the segmentation results obtained by LERF [27] scene representation. It’s essential to note that neither our method nor LERF was initially intended for the segmentation task. Our primary aim is to evaluate the pixel accuracy of the relevance map computed from the rendered CLIP features. We can further post-process and refine our 2D segmentation results by SAM [28] and get the results shown as ‘Ours-Refined’.
@@ 5.2 Unsupervised Segmentation
In following experiments we use FMGS to segment queries and evaluate their segmentation masks. Note that our method is not delicately designed for the segmentation task. We lack a dedicated segmentation header for predicting segmentation masks, nor do we explicitly partition the scene at the object level. We have examined the open-vocabulary language understanding capability of FMGS in the object detection experiments discussed in the above section. Our primary objective for doing this segmentation evaluation is to assess the pixel-level accuracy of the rendered CLIP features obtained from the trained scene representation. Segmentation relies on matching these rendered CLIP features to the embeddings of the provided semantic labels.
@@
Baselines: We conduct a direct comparison of our method with LERF [27]. To perform semantic segmentation, we initially obtain relevancy scores by computing the cosine similarity between the rendered CLIP feature and the embeddings of all class labels (this is different from the relevancy score calculation with auxiliary canonical phrases involved in Sec. 4.3.). These relevancy scores serve as segmentation logits, and we subsequently apply the softmax function to convert them into probabilities. Each pixel is then assigned a semantic class label corresponding to the maximum probability. Note that LERF [27] requires a scale factor when rendering CLIP features, and we report the best segmentation results that can be achieved by LERF by selecting the best scales for each ray. It’s also important to note that both LERF and our method encounter challenges in discerning the semantic labels of backgrounds when presented with visibility-limited close views and lack of context. Therefore, we have replaced the original background labels, including ‘white sheet’, ‘wood wall’, ‘grey sofa’, and ‘lime wall’, with a more general label ‘background’ when testing LERF and our method.
Additionally, for comprehensive reference, we present results obtained using the dedicated 3D-OVS method [33] for the segmentation task. However, it is worth emphasizing that comparing object detection methods like ours and LERF [27] to 3D-OVS is not entirely equitable, as acknowledged in the paper of 3D-OVS [33]. 3D-OVS [33] has prior access to segmentation class labels and distill class-related information into the radiance field during training. In contrast, neither LERF nor our methods have access to class labels during scene representation training. Consequently, the trained 3D-OVS scene representation can only be effectively employed for querying the classes known before training, and does not support arbitrary semantic queries beyond the trained classes. Furthermore, we compare to a 2D ceiling approach [31], OV-Seg, which is directly trained for open-vocabulary semantic segmentation by fine-tuning CLIP on masked image regions and text descriptions. OV-Seg is supervised with mask-category pairs, while ours and LERF are completely unsupervised.
Evaluation Results The segmentation experiment results are presented in Table 2 and Figure 8. Notably, our approach outperforms LERF [27] by a significant margin across all cases. This superior performance can be attributed to the higher quality of our rendered CLIP feature compared to the one produced by LERF. Our method exhibits more concentrated high relevancy around the queried objects, showcasing the advantage of our semantic scene representation, which maintains high semantic consistency in local proximity. We refine our 2D segmentation results using SAM [28] by assigning class labels to SAM segments through a majority voting based on pixel labels obtained from our method. The refined result is displayed in the ‘Ours-Refined’ row of Table 2 and Figure 8.
@@ 5.3.1 Hybrid CLIP feature
In this ablation study, we investigated using a single scale of CLIP features at the finest scale level, rather than our hybrid CLIP features, which are obtained by averaging multiple-scale CLIP features extracted from patches at different resolutions. As demonstrated in Table 3, the hybrid CLIP feature for supervision is greatly important. The scene understanding capability is severely compromised when employing only a single-scale CLIP feature for supervision (denoted as ‘W/O Hybrid CLIP’). The inferior performance observed with the use of a single-scale CLIP feature stems from its potential inadequacy in capturing sufficient contextual information within the image. By contrast, our hybrid CLIP feature encompasses a significantly larger receptive field, enhancing its contextual awareness.
To conduct a comprehensive analysis, we also compare our method of using a single hybrid CLIP supervision (‘Ours’) to the approach of utilizing CLIP supervision at multiple individual scales, as proposed by LERF [27], which we denote as ‘W/ LERF CLIP’. During its training, randomly sampled scale factors are concatenated with the intermediate MHE feature vector $\mathbf{q}$ (refer to Sec. 4.1) to decode CLIP features at each scale. At inference time, the optimal scale factors at each pixel location are determined by an exhaustive search among 30 uniformly distributed scales. The best scale factor is selected based on its relevance score to the query. Consequently, this approach is at least 30 times slower than our method with a single hybrid CLIP supervision during inference. Despite its inefficiency, the accuracy achieved by ‘ W/ LERF CLIP’ ($78.8\%$) significantly lags behind that of ‘Ours’ ($93.2\%$), as shown in Table 3, underscoring the efficacy and superior efficiency of our proposed method. It is worth noting that while it is plausible that ‘W/ LERF CLIP’ could achieve better performance with much more training iterations, this aspect falls beyond the scope of our current investigation.
@@ 5.3.2 Pixel-alignment loss
5.3.3 Scene Representation
To evaluate the effectiveness of our hybrid scene presentation, which integrates 3D Gaussians for geometry and appearance representation alongside MHE for efficient semantic embedding, we compare it with a vanilla scene representation method that solely employs Gaussians without MHE by attaching semantic embeddings to each Gaussian (‘W/O MHE’). Notably, for fair comparisons, we consider only the selected Gaussians with sufficient opacity and 2D radius in the vanilla method (as detailed in Sec. 4.4) identical to our method (‘Ours’). These attached semantic embeddings share the same dimensionality as the intermediate MHE feature vector $\mathbf{q}$ in our method (see Sec. 4.1). They are also decoded by two MLPs, $MLP^{CLIP}_{\phi}$ and $MLP^{DINO}_{\psi}$, to obtain CLIP and DINO features while rendering the feature maps.
As indicated in Table 3, ‘W/O MHE’ achieves an object detection accuracy of 84.4%, which is inferior to ‘Ours’ with its hybrid scene representation. The superior performance of ‘Ours’ can be attributed to its ability to render CLIP feature maps at a higher quality over ‘W/O MHE’. Additionally, in terms of scene representation complexity, the MHE component of our method maintains a constant number of parameters across the five scenes of the LERF dataset, whereas the parameter count of the semantic embeddings in ‘W/O MHE’ varies with the number of Gaussians and increases by 84.1%, 113.5%, 2.14%, 174.2%, and 136.9% compared to ‘Ours’. This substantial memory demand of ‘W/O MHE’ can pose even greater challenges in large-scale scenarios.
@@ 6 Discussion and Limitations
In terms of limitations, FMGS currently relies heavily on the presence of high-quality and calibrated input images, a limitation shared with NeRF-based approaches. Additionally, the performance of FMGS is entirely contingent on the quality of the base foundation models used for training the feature fields. It is conceivable that a model better suited for localizing language within images could yield improved feature field quality. Furthermore, to enhance performance in semantic segmentation tasks, it is advisable to embed a specialized segmentation foundation model, such as SAM [28, 9], into our scene representation. Unlike using SAM solely for post-refinement of our 2D segmentation results, directly distilling SAM results into the 3D space offers the advantage of enforcing multi-view consistency, potentially leading to improved performance.